# New Linear Model of a Composite Energy Storage System with Realizable Dispatch Guarantees

### v1

@@ Abstract
To optimize battery dispatch, a model is required that can predict the state of charge (SOC) trajectory and ensure dispatch is admissible (i.e., does not lead to unexpected SOC saturation). But battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we consider a composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery SOC trajectory while ensuring admissibility of the resulting (aggregated) power schedule and disaggregation to the individual energy storage elements.
@@ I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation]. Grid-scale BESSs are comprised of several battery racks that can be split into several battery modules and further into battery cells. When optimizing BESS dispatch the individual elements are lumped as one composite battery and the composite dispatch is shared equally between the battery elements. Optimizing composite dispatch is a non-convex problem since a complementarity constraint is needed to avoid simultaneous charging and discharging. The complementarity constraint can be formulated with a non-linear constraint, e.g., $P_{\text{c}}[k]P_{\text{d}}[k]=0$, or by the use of binary variables. To obtain a convex formulation, the complementarity constraint is typically omitted. The optimal power trajectories then exhibit simultaneous charging and discharging, but only the net power trajectory is applied. Implementing the net power trajectory results in a discrepancy between the predicted optimal SOC trajectory and the realized SOC trajectory [RobustNawaf, D.Pozo_paper]. This discrepancy can lead to unexpected SOC saturation, resulting in un-realizable power trajectories. In [RobustNawaf, Mazen_CDC_paper], a robust linear battery model is presented as an inner approximation that guarantees that any optimal input power trajectory is realizable, but it is conservative over longer prediction horizons. Previous work has always assumed equal power-sharing between battery elements is enforced. This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by disaggregation to the individual elements without the equal power sharing paradigm. The paper characterizes the set of realizable composite power trajectories that can be disaggregated to the elements while ensuring element-wise power limits, energy limits, and complementarity constraints are satisfied. A priority-based strategy is employed to perform disaggregation.
@@ II Composite Battery Model
A general BESS is shown in Fig 1, where $P_{\text{c/d}}[k]$, $\eta_{\text{c/d}}$, and $E[k]$ are the composite charging/discharging power, charge/discharge efficiency and the SOC at timestep $k$, respectively.
@@
The BESS shown in Fig 1 is a composite of $N$ identical elements. Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery. For element $i\in\{1,\ldots,N\}$, the charge/discharge power and the SOC are denoted by $P^{i}_{\text{c/d}}[k]$ and $E^{i}[k]$, respectively.
A BESS dispatch needs to ensure that power limits, energy limits, and complementarity constraints are satisfied for all elements
@@
$\begin{split}&E^{i}[k+1]=E^{i}[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}^{i}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}[k]\end{split}$ (1a)
$E^{i}[0]=E^{i}_{0}$ (1b)
$0=P_{\text{c}}^{i}[k]P_{\text{d}}^{i}[k]$ (1c)
$0\leq E^{i}[k]\leq{E}_{\text{max}}$ (1d)
$0\leq P_{\text{c/d}}^{i}[k]\leq{P}_{\text{c/d,max}}$ (1e)
@@
$\begin{split}&E[k+1]=E[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
$E[0]=NE_{0}$ (2b)
@@
$\forall i=1,\ldots,N$ and $\forall k=0,\ldots,K-1$, where $\Delta t$ is the time step. Note, it is assumed that the individual elements are identical with the same power and energy limits, ${P}_{\text{c/d,max}}$ and ${E}_{\text{max}}$. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$, as shown in (2). Typically, the composite battery system maintains balanced SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing dispatch across its elements, which means that a feasible composite dispatch implicitly satisfies all elements’ constraints in (1). Under equal power sharing, a single complementarity constraint at the composite battery level ($P_{\text{c}}[k]P_{\text{d}}[k]=0$) is sufficient to satisfy (1c) for all $i$.
However, under unequal power sharing, this does not hold. In fact, with unequal power sharing, we can omit the complementarity constraint at the composite level, but then have to disaggregate the composite power trajectory across the $N$ elements while satisfying (1c). That is, unequal power sharing increases capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints. Thus, this paper presents a priority-based disaggregation strategy that engenders a provably convex set of realizable composite power trajectories.
@@ III Composite Trajectory Disaggregation
A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}[l]\}_{i=1}^{N}$. The controller operates on a faster timescale with $L=MK$ timesteps and $M\geq 1$ and ensures that for each time-step on the fast timescale, $l\in\{0,\ldots,L-1\}$, we have
$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k],$ (3)
where $k=\lfloor\frac{l}{M}\rfloor$. One such controller is the presented priority stack controller discussed below. However, the priority stack controller does not guarantee that constraints in (1) are met, even if the composite power trajectory satisfies the composite power and energy limits. Specifically, under the priority stack policy, (1c) and (1d) may be violated. However, this paper develops sufficient conditions that with the priority stack controller enables characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
@@ III-A Priority Stack Controller (PSC)
Based on the SOC, the priority stack controller (PSC) sorts all $N$ elements every (fast) time-step $l$. The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority. When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack. This process is repeated every time-step $l$, adapting to changing element SOCs. The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
@@
1:Given $P^{\ast}_{\text{c}}[k],P^{\ast}_{\text{d}}[k]$ subject to (2)
@@
4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P^{\ast}_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
@@
Next, we analyze the effects of the PSC on (1), on complementarity conditions (1c), and on energy limits (1d).
@@ III-B Guaranteeing Element Complementarity Constraints
To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously. Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy:
@@
However, considering (4), we can tighten (5) as follows:
@@
The feasible region dictated by (5) and (6) is shown in Fig. 2.
Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right).
@@
If $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy (6) then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (5).
@@
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c/d}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$.
Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (6), we get
Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$, satisfying (5). ∎
Thus, adding (6) to the composite formulation in (2) will satisfy element complementarity constraints under PSC. Similarly, we want to relate element energy limits to those in the composite formulation.
III-C Relating Element and Composite Energy Limit Constraints
Under PSC, the battery elements are charged/discharged individually every $l$, which results in different SOC trajectories. The spread of element SOCs can impact realizability of the composite dispatch especially for larger element SOC spread. This complexity can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits, which tightens the composite energy limits as
$N\Delta E_{\text{max}}\leq E[k]\leq N(E_{\text{max}}-\Delta E_{\text{max}}),$ (8)
where $\Delta E_{\text{max}}$ is given by
$\Delta E_{\text{max}}:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right).$ (9)
Proposition 2.
A composite battery dispatch that satisfies (8), will not violate any individual element’s energy limit under PSC, provided that $\Delta E(0)\leq\Delta E_{\text{max}}$.
Proof.
Given element $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
@@ III-C Relating Element and Composite Energy Limit Constraints
From Property 1 and power limits, we have $-P_{\text{c,max}}\leq(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l))\leq 0$ and $0\leq(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l))\leq P_{\text{d,max}}$, which implies
$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l)$ (11a)
and
$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\Delta E_{\text{max}},$ (11b)
where the last inequality in (11b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Now, we can bound $\Delta E(l+1)\leq\max\left(\Delta E(l),\Delta E_{\text{max}}\right)=\Delta E_{\text{max}}.$ Therefore, $\Delta E(l)\leq\Delta E_{\text{max}}\Rightarrow\Delta E(l+1)\leq\Delta E_{\text{max}}$, which implies that for a composite battery with initial element SOC $E^{i}(0)$, if $\Delta E(0)\leq\Delta E_{\text{max}}$, then $\Delta E(l)\leq\Delta E_{\text{max}}\forall l=0,\ldots,MK$

### v2

# Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees

@@ Abstract
To optimize battery dispatch, a model is required that can predict the state of charge (SOC) trajectory and ensure dispatch is admissible (i.e., does not lead to unexpected SOC saturation). However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we consider a composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery SOC trajectory while ensuring admissibility of the resulting (aggregated) power schedule and disaggregation to the individual energy storage elements.
@@ I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper]. Grid-scale BESSs are comprised of several battery racks that can be split into several battery modules and further into battery cells. When optimizing BESS dispatch the individual elements are lumped as one composite battery and the composite dispatch is shared equally between the battery elements. Optimizing composite dispatch is a non-convex problem since a complementarity constraint is needed to avoid simultaneous charging and discharging. The complementarity constraint can be formulated with a non-linear constraint, or by the use of binary variables. To obtain a convex formulation, the complementarity constraint is typically omitted. The optimal power trajectories then exhibit simultaneous charging and discharging, but only the net power trajectory is applied. Implementing the net power trajectory results in a discrepancy between the predicted optimal SOC trajectory and the realized SOC trajectory [RobustNawaf, D.Pozo_paper]. This discrepancy can lead to unexpected SOC saturation, resulting in un-realizable power trajectories. In [RobustNawaf, Mazen_CDC_paper], a robust linear battery model is presented as an inner approximation that guarantees that any optimal input power trajectory is realizable, but it is conservative over longer prediction horizons. Previous work has always assumed equal power-sharing between battery elements is enforced.
This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements. The paper characterizes the set of realizable composite power trajectories that can be disaggregated to the elements while ensuring element-wise power limits, energy limits, and complementarity constraints are satisfied. A priority-based strategy is employed to perform disaggregation.
@@ II Composite Battery Model
We consider a generic BESS as shown in Fig 1. It is composed of $N$ identical elements. Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery.
@@
For element $i$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$. The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively. A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$). Note that all elements have identical power and energy limits. This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$.
@@
$\begin{split}&E^{i}(l+1)=E^{i}(l)+\\ &\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)\end{split}$ (1a)
$E^{i}(0)=E^{i}_{0}$ (1b)
$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)$ (1c)
$0\leq E^{i}(l+1)\leq{E}_{\text{max}}$ (1d)
$0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1e)
@@
$\begin{split}&E[k+1]=E[k]+\\ &\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
$E[0]=\sum_{i}E^{i}_{0}$ (2b)
@@
The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$. We denote these scheduling time steps with index $[k]$ in square brackets, where $k=0,\ldots,K-1$ and $L=MK$. Constant power in step $k$ implies the following constraint for all $l,k$ with $k=\lfloor\frac{l}{M}\rfloor$:
$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k].$ (3)
Typically, a composite BESS is modeled and operated in a balanced way: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing across its elements. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (2). Under equal power sharing, a single complementarity constraint at the composite battery level, (2c), is sufficient to satisfy (1c) for all $i$ and all $l$. Under balanced operation, a feasible composite dispatch implicitly satisfies all elements’ constraints in (1).
However, under unequal power sharing, this does not hold. In fact, with unequal power sharing, we can omit the complementarity constraint at the composite level, but then have to disaggregate the composite power trajectory across the $N$ elements while satisfying (1c). That is, unequal power sharing increases the capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints. Confronting this challenge, this paper presents a priority-based disaggregation strategy that engenders a provably convex linear set of realizable composite power trajectories. These trajectories permit simultaneous charging and discharging of the BESS and approximate the full solution space arbitrarily closely in well-defined limits.
@@ III Composite Trajectory Disaggregation
A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}(l)\}_{i=1}^{N}$. One such controller is the presented priority stack controller discussed below. Sufficient conditions are developed that – in combination with the controller – enable the characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
@@ III-A Priority Stack Controller (PSC)
Every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC. The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority. When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack. This process is repeated every time-step $l$, adapting to changing element SOCs. The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
@@
1:Given $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$
@@
4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
@@ III-B Guaranteeing Element Complementarity Constraints
Next, we analyze the effects of the PSC on (1), on complementarity conditions (1c), and on energy limits (1d). To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously. Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy:
@@
Proposition 1.
Under the PSC, if $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy
@@
then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (5).
Proof.
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$. Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (6), we get
Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$ for any $l$. ∎
Thus, adding (6) to the composite formulation in (2) will satisfy element complementarity constraints under PSC. The feasible regions dictated by (5) and (6) are shown in Fig. 2.
Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are not considered here.
III-C Relating Element and Composite Energy Limit Constraints
Similarly, we want to relate element energy limits to those in the composite formulation. Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOC trajectories. The spread of element SOCs can impact realizability of the composite dispatch when individual elements hit their SOC limits. Constraint violations can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
@@ III-C Relating Element and Composite Energy Limit Constraints
Under the PSC, a composite dispatch that satisfies
$N\varepsilon\leq E[k]\leq N(E_{\text{max}}-\varepsilon),$ (8)
with
$\varepsilon:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right),$ (9)
will not violate any individual element’s energy limit, provided that $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$.
@@
Define the maximum difference in SOC across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOC spread). Consider element SOCs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
@@
From Property 1 and (1e), we have $-P_{\text{c,max}}\leq P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\leq 0$ and $0\leq P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\leq P_{\text{d,max}}$, which implies
$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l),$ (11a)
$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\varepsilon,$ (11b)
where the inequality in (11b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$. Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$, we maintain $\Delta E(l)\leq\varepsilon$ for all $l$.
Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (8) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$. Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOC. Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$. Combined, these inequalities show that (1d) holds for all $i$ and all $l$. ∎